Sametion of a^3 cos (B-C) find the answer related to properties of triangle

Grade:12th pass

1 Answers

Arun
5 years ago
X = a³cos(B-C) + b³cos(C-A) +c³cos(A-B) 
 
From sine rule and 3A trig identity a³ = 8R³sin³(A) = 2R³(3sin(A)−sin(3A)) 
 
So X = 2R³ ∑( 3sin(A)−sin(3A) )cos(B−C) = R³Y where ∑ is over cycles 
 
2sin(A)cos(B−C) = sin(A+B−C) + sin(A−B+C) = sin(2C) + sin(2B) 
 
2sin(3A)cos(B−C) = sin(3A+B−C) + sin(3A−B+C) 
 
with similar expressions for B,C,A & C,A,B 
 
∴ Y = 3∑sin(2C) + 3∑sin(2B) − ∑sin(3A+B−C) − ∑sin(3A−B+C) 
 
∑sin(3A+B−C) + ∑sin(3A−B+C) = sin(3A+B−C) + sin(3B+C−A) + sin(3C+A−B) + 
……………………………………... .sin(3A−B+C) + sin(3B−C+A) + sin(3C−A+B) … (i) 
 
But (3A+B−C) + (3C−A+B) = 2(A+B+C) = 2π so sin(3A+B−C) + sin(3C−A+B) = 0 with similar for other pairs of terms in (i). So (i) is zero. 
 
∴ Y = 6∑sin(2A) = 24sin(A)sin(B)sin(C) (an easily proved identity for a triangle) 
 
Y = 24abc/(8R³) = 3abc/R³ → X = 3abc